# A Distributionally-Robust Framework for Nuisance in Causal Effect Estimation

### v1

@@ 5.1 Settings
We basically followed the setup of [10] for synthetic data generation processes and model hyperparameters. We generated $d=25$ dimensional multivariate normal covariates including 5 confounding factors that affect outcome and treatment selection, 5 outcome-related factors, and 5 purely CATE-related factors. The true outcome and CATE models were nonlinear. Other details are described in Appendix 0.D.
@@
Most well-established semi-synthetic datasets have real-world covariates and synthesized outcomes and do not reproduce up to noise heterogeneity, while the Twins dataset from [1] had both potential outcomes recorded. The first-year mortality of twins born at low birth weights was treated as potential outcomes for the heavier and lighter-born twins, respectively, and weighted subsampling was performed to reproduce the missing counterfactuals. The test target is the difference between the noisy potential outcomes $\tau^{(n)}=y_{1}^{(n)}-y_{0}^{(n)}$ instead of CATE $\mathbb{E}[\tau^{(n)}|x]$. We followed [11] for sampling strategy and other details.
Also, the Jobs dataset from [23] has a randomized test set based on an experimental job training program and an observational training set. Although we do not have both potential outcomes, we can substitute the true CATE label in the PEHE with the transformed outcome with known true propensity in the randomized test set, as proposed in [12]. A detailed justification for this evaluation is explained in Appendix 0.B.
@@
We compared several representative methods for causal inference with DNN. TNet was a simple plug-in method that estimates each potential outcome with two separate networks and then outputs the difference between their predictions. SNet and SNet3 were decomposed representation-based methods that shared representation extractors for outcome and propensity estimation layers. They have three kinds of extractors, namely, outcome-only, propensity-only, and shared representation for confounders. SNet3 was a modified version, not using weighting in [10], of what was originally proposed as DR-CFR [15] and DeR-CFR [41] for simultaneous optimization of the weights and the outcome model. SNet3 with reweighting is presented in Appendix 0.E. SNet was an even more flexible architecture than SNet3 that had shared and independent extractors for each potential outcome, proposed in [10]. DRNet was an DNN implementation of DR-Learner [21], in which networks in the first step were independent for each potential outcome and propensity. In the second step, a single CATE network was trained.
@@
We set the candidates of the hyperparameters as $\alpha_{0}\in\{1,10\}$, $\gamma\in\{1.5,2,3\}$, and $\beta\in\{10,100,300\}$. For the experiment with Twins, we fixed them as $\alpha_{0}=10,\gamma=1.5,\beta=100.$ Model selection, including early stopping, was performed on the evidence measure of each method with $r=30$% of the training data held out for validation and the pre-trained weights are used for the proposed methods.
@@ 5.2 Results
The results are shown in Table 1 and Table 2. Our proposed NuDRNet and NuSNet outperformed or was at least comparable to their baseline methods, DRNet and SNet, respectively. On the other hand, representation-based methods (SNet3 and SNet) outperformed the transformed outcome methods (DRNet and NuDRNet). The shared representation extractor of the confounding factors could be an effective inductive bias, especially with small sample sizes. SNet is overall more accurate than SNet3 since it can also share parameters for components common to both potential outcomes. Here, the comparison with DRNet-o using true propensity is interesting. DRNet uses estimated propensity while DRNet-o uses the true one, and DRNet-o showed superior performance in the most cases. NuDRNet, despite being propensity-agnostic, recovered an accuracy comparable to DRNet-o, which illustrates the benefit of robustness against the ambiguity of nuisance estimation.
Table 3: MSE for noisy CATE on real-world datasets (mean $\pm$ standard error over 5 runs). It’s important to note that for real-world datasets, we can only observe a noisy alternative of the true label $\tau$. That is, even as the sample size $N$ approaches infinity, the MSE for this noisy CATE estimate does not converge to zero. Only comparisons between methods are meaningful; comparisons to zero are not applicable.
@@
Table 1 also shows the results in the multiplicative noise setting. NuDRNet outperformed other baselines when the sample size was relatively sufficient. The pessimistic evaluation with more emphasis on hard instances would be a reasonable explanation for the superiority of the proposed method. Even though representation decomposition should also be useful in the MN setting since the data generation model was the same as the AN setting except for noise, the weighting approach was superior to the representation decomposition method without weighting. Again, NuDRNet recovered the DRNet-o with the true propensity.
@@
Experiments on Twins data also showed the superiority of NuDRNet and NuSNet in most cases as in Table 3. Note that the test target $\tau^{(n)}$ is noisy, and the value contains the noise variance. Table 3 also showed the results on Jobs data, which exhibits similar trends. Note that the evaluation metric for Jobs, the MSE with respect to the transformed outcome, also contains constant noise. Although it may seem that there is not much difference between any of the methods compared to zero, it is possible that there is a significant improvement in PEHE if unobservable constant noise was removed.
@@ Appendix 0.A Proofs
where $\sigma_{n}\sim\{\pm 1\}$ denotes the Rademacher random variable with equal probability of being 1 or -1, i.e., $P(\sigma_{n}=1)=P(\sigma_{n}=-1)=\frac{1}{2}$. We applied Jensen’s inequality in 28 and the element-wise contraction lemma of the Rademacher complexity  in 32:
($\mathfrak{R}(\{\ell_{n}(a_{n})\})\leq\mathfrak{R}(\{\rho_{n}a_{n}\}),$ (34)
@@ Appendix 0.B Equivalence between the transformed outcome and the CATE in expectation
Our empirical risk and the evaluation criteria are based on the method of transformed outcome . The IPW transformed outcome $z$ in LABEL:eq:_transformed_outcome_PW is equivalent to the CATE in the sense of its conditional expectation:
@@
The MSE on $z$ is equivalent to our final metric PEHE except for a constant term. The same equivalence can be derived for the doubly robust transformed outcome in 14. This equivalence justifies our employed MSE on $z$ as the empirical risk $\hat{L}$ and the evaluation metric.
@@ Appendix 0.C Algorithm for NuSNet
Algorithm 2 presents the full algorithm of our proposed method in Section 4.3.
@@
1:  Train $f_{0},f_{1},\pi$ by an arbitrary supervised learning method, e.g.: $\hat{f}_{a}\leftarrow\operatorname*{arg\,min}_{f_{a}}\frac{1}{N}\sum_{n:a^{(n)}=a}(y^{(n)}-f_{a}(x^{(n)}))^{2}$ for each $a\in\{0,1\}$,$\pi_{0}\leftarrow\operatorname*{arg\,min}_{\pi}-\frac{1}{N}\sum_{n}a^{(n)}\log\pi(x^{(n)})+(1-a^{(n)})\log(1-\pi(x^{(n)}))$
@@
6:        Update parameters with objective 19 and step sizes $\eta_{\theta_{0}},\eta_{\theta_{1}},\eta_{\phi}$ from optimizes:
@@ 0.D.1 Simulation environment
[…]
@@
where $\sigma$ was the sigmoid, $\xi$ was the strength of selection and was set as $\xi=3$, and $\omega$ was adaptively set so that the median of the inside $\sigma$ would be 00.
@@ Appendix 0.E Additional results
The results are shown in Table 4 and Table 5. In our tested settings, these methods did not perform well. Despite the asymptotic guarantees of these methods, they tend to set extreme weights, making it difficult to achieve consistent performance with realistic sample sizes, especially when models are complex or the overlap is limited.

### v2

@@ 5.1 Settings
We basically followed the setup of [curth2021nonparametric] for synthetic data generation processes and model hyperparameters. We generated $d=25$ dimensional multivariate normal covariates including 5 confounding factors that affect outcome and treatment selection, 5 outcome-related factors, and 5 purely CATE-related factors. The true outcome and CATE models were nonlinear. Other details are described in Appendix˜0.D.
@@
Most well-established semi-synthetic datasets have real-world covariates and synthesized outcomes and do not reproduce up to noise heterogeneity, while the Twins dataset from [almond2005costs] had both potential outcomes recorded. The first-year mortality of twins born at low birth weights was treated as potential outcomes for the heavier and lighter-born twins, respectively, and weighted subsampling was performed to reproduce the missing counterfactuals. The test target is the difference between the noisy potential outcomes $\tau^{(n)}=y_{1}^{(n)}-y_{0}^{(n)}$ instead of CATE $\mathbb{E}[\tau^{(n)}|x]$. We followed [curth2021inductive] for sampling strategy and other details.
Also, the Jobs dataset from [lalonde1986evaluating] has a randomized test set based on an experimental job training program and an observational training set. Although we do not have both potential outcomes, we can substitute the true CATE label in the PEHE with the transformed outcome with known true propensity in the randomized test set, as proposed in [Curth2021ReallyDG]. A detailed justification for this evaluation is explained in Appendix˜0.B.
@@
We compared several representative methods for causal inference with DNN. TNet was a simple plug-in method that estimates each potential outcome with two separate networks and then outputs the difference between their predictions. SNet and SNet3 were decomposed representation-based methods that shared representation extractors for outcome and propensity estimation layers. They have three kinds of extractors, namely, outcome-only, propensity-only, and shared representation for confounders. SNet3 was a modified version, not using weighting in [curth2021nonparametric], of what was originally proposed as DR-CFR [hassanpour2019counterfactual] and DeR-CFR [wu2022learning] for simultaneous optimization of the weights and the outcome model. SNet3 with reweighting is presented in Appendix˜0.E. SNet was an even more flexible architecture than SNet3 that had shared and independent extractors for each potential outcome, proposed in [curth2021nonparametric]. DRNet was an DNN implementation of DR-Learner [kennedy2020towards], in which networks in the first step were independent for each potential outcome and propensity. In the second step, a single CATE network was trained.
@@
We set the candidates of the hyperparameters as $\alpha_{0}\in\{1,10\}$, $\gamma\in\{1.5,2,3\}$, and $\beta\in\{10,100,300\}$. For the experiment with Twins, we fixed them as $\alpha_{0}=10,\gamma=1.5,\beta=100.$ Model selection, including early stopping, was performed on the evidence measure of each method with r=30r=30% of the training data held out for validation and the pre-trained weights are used for the proposed methods.
@@ 5.2 Results
The results are shown in Table˜1 and Table˜2. Our proposed NuDRNet and NuSNet outperformed or was at least comparable to their baseline methods, DRNet and SNet, respectively. On the other hand, representation-based methods (SNet3 and SNet) outperformed the transformed outcome methods (DRNet and NuDRNet). The shared representation extractor of the confounding factors could be an effective inductive bias, especially with small sample sizes. SNet is overall more accurate than SNet3 since it can also share parameters for components common to both potential outcomes. Here, the comparison with DRNet-o using true propensity is interesting. DRNet uses estimated propensity while DRNet-o uses the true one, and DRNet-o showed superior performance in the most cases. NuDRNet, despite being propensity-agnostic, recovered an accuracy comparable to DRNet-o, which illustrates the benefit of robustness against the ambiguity of nuisance estimation.
Table 3: MSE for noisy CATE on real-world datasets (mean $\pm$ standard error over 5 runs). It’s important to note that for real-world datasets, we can only observe a noisy alternative of the true label $\tau$. That is, even as the sample size NN approaches infinity, the MSE for this noisy CATE estimate does not converge to zero. Only comparisons between methods are meaningful; comparisons to zero are not applicable.
@@
Table˜1 also shows the results in the multiplicative noise setting. NuDRNet outperformed other baselines when the sample size was relatively sufficient. The pessimistic evaluation with more emphasis on hard instances would be a reasonable explanation for the superiority of the proposed method. Even though representation decomposition should also be useful in the MN setting since the data generation model was the same as the AN setting except for noise, the weighting approach was superior to the representation decomposition method without weighting. Again, NuDRNet recovered the DRNet-o with the true propensity.
@@
Experiments on Twins data also showed the superiority of NuDRNet and NuSNet in most cases as in Table˜3. Note that the test target $\tau^{(n)}$ is noisy, and the value contains the noise variance. Table˜3 also showed the results on Jobs data, which exhibits similar trends. Note that the evaluation metric for Jobs, the MSE with respect to the transformed outcome, also contains constant noise. Although it may seem that there is not much difference between any of the methods compared to zero, it is possible that there is a significant improvement in PEHE if unobservable constant noise was removed.
@@ 6 Conclusion
Acknowledgements
The Version of Record of this contribution is published in the Neural Information Processing, ICONIP 2025 Proceedings and is available online at https://doi.org/10.1007/978-981-95-4094-5_19.
@@ Appendix 0.A Proofs
where $\sigma_{n}\sim\{\pm 1\}$ denotes the Rademacher random variable with equal probability of being 1 or -1, i.e., $P(\sigma_{n}=1)=P(\sigma_{n}=-1)=\frac{1}{2}$. We applied Jensen’s inequality in ˜28 and the element-wise contraction lemma of the Rademacher complexity  in ˜32:
$\mathfrak{R}(\{\ell_{n}(a_{n})\})\leq\mathfrak{R}(\{\rho_{n}a_{n}\}),$ (34)
@@ Appendix 0.B Equivalence between the transformed outcome and the CATE in expectation
Our empirical risk and the evaluation criteria are based on the method of transformed outcome . The inverse-probability-weighted (IPW) transformed outcome is
$z^{(n)}=\frac{a^{(n)}y^{(n)}}{\mu(x^{(n)})}-\frac{\left(1-a^{(n)}\right)y^{(n)}}{1-\mu(x^{(n)})}.$ (41)
The IPW transformed outcome $z$ in ˜41 is equivalent to the CATE in the sense of its conditional expectation:
@@
The MSE on $z$ is equivalent to our final metric PEHE except for a constant term. The same equivalence can be derived for the doubly robust transformed outcome in ˜14. This equivalence justifies our employed MSE on $z$ as the empirical risk $\hat{L}$ and the evaluation metric.
@@ Appendix 0.C Algorithm for NuSNet
Algorithm˜2 presents the full algorithm of our proposed method in Section˜4.3.
@@
1: Train $f_{0},f_{1},\pi$ by an arbitrary supervised learning method, e.g.: $\hat{f}_{a}\leftarrow\operatorname*{arg\,min}_{f_{a}}\frac{1}{N}\sum_{n:a^{(n)}=a}(y^{(n)}-f_{a}(x^{(n)}))^{2}$ for each $a\in\{0,1\}$$\pi_{0}\leftarrow\operatorname*{arg\,min}_{\pi}-\frac{1}{N}\sum_{n}a^{(n)}\log\pi(x^{(n)})+(1-a^{(n)})\log(1-\pi(x^{(n)}))$
@@
6:   Update parameters with objective ˜19 and step sizes $\eta_{\theta_{0}},\eta_{\theta_{1}},\eta_{\phi}$ from optimizes:
@@ 0.D.1 Simulation environment
Squaring works on an element-by-element basis where $\mathbf{1}=[1,\cdots,1]^{\top},$ squaring ${\cdot}^{2}$ is element-wise, and the treatment effect $\tau(x)$ was defined as
@@
where $\sigma$ was the sigmoid, $\xi$ was the strength of selection and was set as $\xi=3$, and $\omega$ was adaptively set so that the median of the inside $\sigma$ would be 0.
@@ Appendix 0.E Additional results
The results are shown in Table˜4 and Table˜5. In our tested settings, these methods did not perform well. Despite the asymptotic guarantees of these methods, they tend to set extreme weights, making it difficult to achieve consistent performance with realistic sample sizes, especially when models are complex or the overlap is limited.